OpenStudy (anonymous):

express this in radical form and simplify. 729 1/4

OpenStudy (anonymous):

Do you mean\[729^{\frac{1}{4}}\] Because then remember that \[x^{\frac{1}{n}} = \sqrt[n]{x}\]

OpenStudy (anonymous):

yes i do